OpenStudy (anonymous):

whats the answer? 78 = -2(m + 3) + m

OpenStudy (anonymous):

\[78=-2m-6+m ...... 78=-m-6...........78+6=-m.........84=-m....m=-84\]

OpenStudy (anonymous):

so the answer is 84? the ending got cut off!

OpenStudy (anonymous):

-84

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

welcome